# Large Language Model as a Policy Teacher for Training Reinforcement Learning Agents

### v5

@@ 1 Introduction
Large Language Models (LLMs) have revolutionized the field of artificial intelligence. These models are trained with an internet-scale text corpus, enabling them to exhibit remarkable capabilities such as natural language generation, question answering, and translation (Brown et al., 2020; Du et al., 2022; Chiang et al., 2023). Previous work suggests that these models contain vast general knowledge about the world and are capable of solving complex reasoning problems (Radford et al., 2019; Brown et al., 2020; Wei et al., 2022). Recently, several works have attempted to use LLMs to generate action plans in an embodied environment (Ahn et al., 2022; Wang et al., 2023a; Driess et al., 2023; Song et al., 2023; Sha et al., 2023; Mao et al., 2023). However, LLMs face challenges in generating effective end-to-end instructions for specific embodied tasks, especially in real-world dynamic scenarios. This limitation arises from two key factors. Firstly, LLMs do not possess the appropriate task incentives during the training process. Secondly, these models lack the capability to actively interact with the environment and gather real-time data (Carta et al., 2023).
Furthermore, the utilization of LLMs often requires substantial computational resources, e.g., memory and power. These requirements render their deployment impractical and expensive, especially when considering their use on lightweight edge devices. These challenges motivate us to address the following question:
@@
  • We propose LLM4Teach, a policy distillation approach to address the limitations of LLM and RL-based agents for making sequential decisions in embodied settings.
@@
  • As a byproduct, we demonstrate that relying solely on LLM can result in various types of incorrect decisions in embodied settings, while LLM4Teach offers an effective approach to mitigate or avoid the influence caused by such incorrect decisions. We also verify that offering uncertainty-aware guidance rather than deterministic guidance through LLM can improve the sample efficiency for the student agent.
@@ 2 Related Work
In this paper, we consider an algorithmic agent operating in a dynamic environment. The agent is required to make a series of decisions and take actions based on the current state of the environment, employing a specific policy to successfully complete a designated task. Here we provide a brief overview of relevant research in the literature.
@@ 2.1 LLM-based Agents
In contrast to the previously mentioned methods, our approach involves utilizing the LLM solely during the training phase to distill task-specific knowledge from the LLM into a RL-based agent. Subsequently, during the testing phase, only the lightweight student agent operates independently without dependence on the LLM.
@@ 2.2 LLM Assisted RL
Several studies have investigated the potential of utilizing LLMs to support the standard RL process by tapping into the general knowledge embedded in LLMs. For example, Kwon et al. (2023); Yu et al. (2023) and Klissarov et al. (2023) address complex scenarios where defining desired behaviors using a simple reward function is challenging. These studies employ LLMs to assist in assigning rewards. Kwon et al. (2023) use LLMs as proxy reward functions to automatically label trajectory data with rewards, while Yu et al. (2023) utilize LLMs to flexibly define reward parameters for optimizing and completing various robot tasks. In a different approach, Klissarov et al. (2023) leverage an offline dataset of behaviors and use LLMs’ preferences over pairs of randomly sampled trajectories to construct a reward model. Furthermore, Du et al. (2023) and Colas et al. (2023) focus on learning diverse behaviors without relying on reward supervision, employing LLMs to generate novel goals during exploration in the environment.
In contrast to these previous works, our approach focuses on leveraging prior knowledge about the target task to enhance the initial exploration stage of an RL agent. This allows us to train the policy model with significantly less data, thereby improving the efficiency of the learning process.
@@ 2.3 Learning from Teacher Agents
Prior research has sought to improve the inefficiencies of tabula rasa RL by utilizing existing teacher agents to guide the learning process of a specialized student agent for the specific problem at hand (Da Silva et al., 2020; Agarwal et al., 2022). These instructions can manifest as demonstrations (Schaal, 1996), scalar feedback (Knox and Stone, 2009), or action advice (Da Silva et al., 2017). Jump-start RL involves the use of a teacher agent to assist in gathering high-quality data during the initial exploration phase of RL (Uchendu et al., 2023). Kickstarting RL combines on-policy distillation with RL, prompting the student agent to emulate the teacher’s behavior while optimizing for accumulated returns (Schmitt et al., 2018). Matthews et al. (2022) further extends this approach to hierarchical policies, transferring pre-trained low-level skill policies as teachers and training the student agent alongside a policy-over-teachers from scratch, which weighs the advice from each teacher agent at every time step.
In contrast, our approach does not rely on specialized teacher agents for the target problem. Instead, we harness the extensive general knowledge embedded in LLMs and pre-trained fundamental skills to construct an LLM-based teacher agent and expedite the learning process of the student agent through on-policy distillation.
@@ 3.3.1 Generating Uncertainty-aware Instructions Using LLM
The process of the student agent learning policies from the teacher agent can be seen as distilling important task-related knowledge from the LLM agent. As demonstrated in Hinton et al. (2015), incorporating uncertainty into knowledge distillation can improve sample efficiency and prevent model over-fitting. Consequently, we propose having the LLM offer uncertainty-aware soft instructions to the student agent. When the student agent sends a text description $c(s)$ to the teacher agent, the teacher agent responds by providing a soft decision $\pi_{T}(\cdot|s)$, i.e., a distribution over available policies, in the following way:
@@ 3.4 On the Learning Process of the Student Agent
where $\mathcal{L}_{\text{RL}}(\theta)$ denotes the traditional loss used in RL algorithms to encode the feedback from the environment. This loss is typically designed to maximize the expected return or rewards obtained by the agent. We incorporate the teacher agent’s guidance into the student agent’s learning process by introducing regular terms $\mathcal{H}\left(\pi_{T}(\cdot|s)||\pi_{\theta}(\cdot|s)\right)$ that describe the differences between teacher and student policies. This term captures the Kullback-Leibler (KL) divergence or Wasserstein distance between the policy of the student agent and the policy $\pi_{T}(\cdot|s)$ of the teacher agent. To control the extent to which the student agent relies on the teacher agent, we introduce an annealing parameter $\lambda$. When $\lambda$ is set to zero, the learning process reduces to a standard RL process without any influence from the teacher agent.
@@ 4.0.1 Simulation Platforms
MiniGrid  offers a customizable grid world environment with various sizes, object types, and objectives, making it a simple representation of grid-based tasks (Chevalier-Boisvert et al., 2023). These tasks pose a challenge for RL methods because of their sparse rewards.
Habitat  is a simulation platform specifically created to support the development of embodied AI systems (Szot et al., 2021). It provides a comprehensive framework for defining and executing various embodied AI tasks, such as navigation, object rearrangement, and question-answering. Additionally, Habitat enables detailed configuration of embodied agents, including their physical attributes and sensor specifications.
@@ 4.0.2 Baseline Methods
LLM soly  which operates in two stages. First, the LLM selects an option from a set of available pre-trained options. Then, a low-level action is generated following the selected option policy. In this configuration, only an LLM-based agent is utilized to make real-time decisions, without the involvement of the student agent. This approach allows us to investigate the potential of our proposed LLM4Teach framework in enabling the student agent to outperform its teacher in completing the desired task.
Hierarchical RL , a RL baseline approach that involves training the student agent with pre-trained option policies (Matthews et al., 2022). In light of the hierarchical nature of the tasks, we explore such hierarchical RL approach in our experiments, so that we can assess the benefits of knowledge distillation from a pre-trained LLM that captures world knowledge.
Baseline RL , a Tabula rasa RL trained from scratch using the proximal policy optimization (PPO) algorithm (Schulman et al., 2017). The policy model structure and the training loss function are set the same as our student agent in LLM4Teach.
@@ 4.1.1 Experimental Setting
We created 4 procedurally generated tasks in the MiniGrid environment (Chevalier-Boisvert et al., 2023): {SimpleDoorKey, ColoredDoorKey, LavaDoorKey and DivergedDoorKey}. In each task, the agents are situated in rooms with varying layouts and their goal is to unlock the exit door using the correct key. In SimpleDoorKey, the agent must explore the room, find a key, and use it to unlock the exit door. In ColoredDoorKey, the exit door can only be unlocked with a key that matches its color, adding complexity for the agent to understand task-specific rules. LavaDoorKey introduces hazard grids (Lava) to the room, requiring the agent to quickly adapt to new elements. DivergedDoorKey presents two exit doors instead of one, allowing the agent to choose either door to escape, emphasizing the importance of using uncertainty-aware instructions to improve overall sample efficiency.
@@ 4.1.3 Ablation Study on Uncertainty-aware Instructions
Figure 4: Ablation study on uncertainty-aware instructions. It shows that the utilization of two type of uncertainty-aware instructions by the teacher agent results in improved sample efficiency for the student agent.
@@
The result of the ablation study is shown in Figure 4. It can be observed that utilizing uncertainty-aware instructions improves the overall sample efficiency compared to using deterministic ones. Moreover, there is no significant disparity in performance between the two approaches for generating uncertainty-aware instructions. The first approach is simpler to implement in practical scenarios but consumes more computational resources due to multiple queries to LLMs, particularly when the observation space is large. On the other hand, the second approach necessitates access to logits, making it applicable only to open-source LLMs.
@@ 4.2.1 Experimental Setting
We select the Vicuna-7b model (Chiang et al., 2023) as the LLM used in LLM4Teach, following a similar prompt design as in previous experiments on Minigrid. Moreover, we utilize visual observations captured by the on-board camera as input queries for the LLM. To enable the LLM-based teacher agent to comprehend these visual inputs, we utilize a preconfigured translator that generates natural language descriptions listing the objects identified in the visual inputs. Alternatively, pre-trained visual-language models such as CLIP (Radford et al., 2021) can also be utilized for this purpose.
@@ 5 Concluding Remarks
Both RL and LLMs have limitations in handling complex sequential decision-making problems. RL often lacks sample efficiency and incurs high exploration costs, while LLMs are prone to decision errors and have high deployment costs. Combining LLMs with RL to overcome these limitations is a natural idea, but creating an effective interface between them poses challenges. LLMs utilize texts as input and output, making them suitable for providing high-level instructions, whereas RL operates at a lower level and uses numerical vectors instead of texts.
Here we present LLM4Teach, a novel framework that combines LLMs and RL for embodied sequential decision-making tasks. Our approach leverages the reasoning capabilities of LLMs to develop a highly capable RL-based student agent. In particular, we use the LLM to provide high-level suggestions on available options for policy training of the student agent. Extensive experiments demonstrate that our student agent outperforms all RL baselines in sample efficiency. Meanwhile, it achieves superior performance to LLM soly in terms of task completion success rate with much fewer computational resources during online testing. For instance, in MiniGrid experiments, the student agent’s model size is 24K compared to LLM’s 130B. Similarly, in Habitat experiments, the student agent’s model size is 10M while LLM’s is 7B.

### v6

@@ 1 Introduction
Large Language Models (LLMs) have revolutionized the field of artificial intelligence. These models are trained with an internet-scale text corpus, enabling them to exhibit remarkable capabilities such as natural language generation, question answering, and translation (Brown et al., 2020; Du et al., 2022; Chiang et al., 2023). Previous work suggests that these models contain vast general knowledge about the world and are capable of solving complex reasoning problems (Radford et al., 2019; Brown et al., 2020; Wei et al., 2022). Recently, several works have attempted to use LLMs to generate action plans in an embodied environment (Ahn et al., 2022; Wang et al., 2023a; Driess et al., 2023; Song et al., 2023; Sha et al., 2023; Mao et al., 2023). However, LLMs face challenges in generating effective end-to-end instructions for specific embodied tasks, especially in real-world dynamic scenarios. This limitation arises from two key factors. Firstly, LLMs do not possess the appropriate task incentives during the training process. Secondly, these models lack the capability to actively interact with the environment and gather real-time data (Carta et al., 2023). Furthermore, the utilization of LLMs often requires substantial computational resources, e.g., memory and power. These requirements render their deployment impractical and expensive, especially when considering their use on lightweight edge devices. These challenges motivate us to address the following question:
@@
  • We propose LLM4Teach, a policy distillation approach to address the limitations of LLM and RL-based agents for embodied sequential decision making.
@@
  • As a byproduct, we demonstrate that relying solely on LLM can result in various types of incorrect decisions in embodied settings, while LLM4Teach offers an effective approach to mitigate or avoid the influence caused by such incorrect decisions. We also verify that offering uncertainty-aware rather than deterministic guidance through LLM can improve the sample efficiency for the student agent.
@@ 2 Related Work
In this paper, we consider an algorithmic agent operating in an open dynamic environment. The agent is required to make a series of decisions and take actions based on the current state of the environment, employing a specific policy to successfully complete a designated task. Here we provide a brief overview of relevant research in the literature.
@@ 2.1 LLM-based Agents
In contrast, our approach involves utilizing the LLM solely during the training phase to distill task-specific knowledge from the LLM into a RL-based agent. Subsequently, during the testing phase, only the lightweight student agent is deployed, which works independently without dependence on the LLM.
@@ 2.2 LLM Assisted RL
Several studies have investigated the potential of utilizing LLMs to support the standard RL process by tapping into the general knowledge embedded in LLMs. For example, Kwon et al. (2023); Yu et al. (2023) and Klissarov et al. (2023) employ LLMs to assist in assigning rewards. Kwon et al. (2023) use LLMs as proxy reward functions to automatically label trajectory data with rewards, while Yu et al. (2023) utilize LLMs to flexibly define reward parameters for optimizing and completing various robot tasks. In a different approach, Klissarov et al. (2023) leverage an offline dataset of behaviors and use LLMs’ preferences over pairs of randomly sampled trajectories to construct a reward model. Furthermore, Du et al. (2023) and Colas et al. (2023) focus on learning diverse behaviors without relying on reward supervision, employing LLMs to generate novel goals during exploration in the environment. In contrast to these previous works, our approach focuses on leveraging prior knowledge about the target task to enhance the initial exploration stage of an RL agent. This allows us to train the policy model with significantly less data, thereby improving the efficiency of the learning process.
@@ 2.3 Learning from Teacher Agents
Prior research has sought to improve the inefficiencies of tabula rasa RL by utilizing existing teacher agents to guide the learning process of a specialized RL agent for a specific target problem (Da Silva et al., 2020; Agarwal et al., 2022). These instructions can manifest as demonstrations (Schaal, 1996), scalar feedback (Knox and Stone, 2009), or action advice (Da Silva et al., 2017). Jump-start RL involves the use of a teacher agent to assist in gathering high-quality data during the initial exploration phase of RL (Uchendu et al., 2023). Kickstarting RL combines on-policy distillation with RL, prompting the student agent to emulate the teacher’s behavior while optimizing for accumulated returns (Schmitt et al., 2018). Matthews et al. (2022) further extends this approach to hierarchical policies, transferring pre-trained low-level skill policies as teachers and training the student agent alongside a policy-over-teachers from scratch, which weighs the advice from each teacher agent at every time step.
In contrast, our approach does not rely on specialized teacher agents for the target problem. Instead, we harness the extensive general knowledge embedded in LLMs to expedite the learning process of the student agent through on-policy distillation for combining pre-trained fundamental skills to complete the target task.
@@ 3.2.1 Generating Uncertainty-aware Instructions Using LLM
The process of the student agent learning policies from the teacher agent can be seen as distilling important task-related knowledge from the LLM agent. As demonstrated in Hinton et al. (2015), incorporating uncertainty into knowledge distillation can improve sample efficiency and prevent model over-fitting. Consequently, we propose having the LLM offer uncertainty-aware soft instructions to the student agent. When the student agent sends a text description $c(s)$ to the teacher agent, the teacher agent responds by providing a soft decision $\pi_{T}(\cdot|s)$, i.e., a distribution over available options, in the following way:
@@ 3.3 On the Learning Process of the Student Agent
where $\mathcal{L}_{\text{RL}}(\theta)$ denotes the traditional loss used in RL algorithms to encode the feedback from the environment. This loss is typically designed to maximize the expected return or rewards obtained by the agent. We incorporate the teacher agent’s guidance into the student agent’s learning process by introducing the regularization term $\mathcal{H}\left(\pi_{T}(\cdot|s)||\pi_{\theta}(\cdot|s)\right)$ that describes the difference between the teacher policy $\pi_{T}(\cdot|s)$ and the student policy. This term captures the Kullback-Leibler (KL) divergence or Wasserstein distance between those two policies. To control the extent to which the student agent relies on the teacher agent, we introduce an annealing parameter $\lambda$. When $\lambda$ is set to zero, the learning process of the student agent reduces to a standard RL process without any influence from the teacher agent.
@@ 4.0.1 Simulation Platforms
MiniGrid
offers a customizable grid world environment with various sizes, object types, and objectives, making it a simple representation of grid-based tasks (Chevalier-Boisvert et al., 2023). These tasks pose a challenge for RL methods because of their sparse rewards.
Habitat
is a simulation platform specifically created to support the development of embodied AI systems (Szot et al., 2021). It provides a comprehensive framework for defining and executing various embodied AI tasks, such as navigation, object rearrangement, and question-answering. Additionally, Habitat enables detailed configuration of embodied agents, including their physical attributes and sensor specifications.
@@ 4.0.2 Baseline Methods
LLM soly
operates in two stages. First, the LLM selects an option from a set of available pre-trained options. Then, a low-level action is generated following the selected option policy. In this configuration, only an LLM-based agent is utilized to make real-time decisions, without the involvement of the student agent. This approach allows us to investigate the potential of our proposed LLM4Teach framework in enabling the student agent to outperform its teacher in completing the desired task.
Hierarchical RL
is an RL baseline approach that involves training the student agent with pre-trained option policies (Matthews et al., 2022). In light of the hierarchical nature of the tasks, we explore such hierarchical RL approach in our experiments, so that we can assess the benefits of knowledge distillation from a pre-trained LLM that captures world knowledge.
Baseline RL
is a Tabula rasa RL baseline trained from scratch using the proximal policy optimization (PPO) algorithm (Schulman et al., 2017). The policy model structure and the training loss function are set the same as our student agent in LLM4Teach.
@@ 4.1.1 Experimental Setting
We created 4 procedurally generated tasks in the MiniGrid environment: {SimpleDoorKey, ColoredDoorKey, LavaDoorKey and DivergedDoorKey}. In each task, the agents are situated in rooms with varying layouts and their goal is to unlock the exit door using the correct key. In SimpleDoorKey, the agent must explore the room, find a key, and use it to unlock the exit door. In ColoredDoorKey, the exit door can only be unlocked with a key that matches its color, adding complexity for the agent to understand task-specific rules. LavaDoorKey introduces hazard grids (Lava) to the room, requiring the agent to quickly adapt to new elements. DivergedDoorKey presents two exit doors instead of one, allowing the agent to choose either door to escape, emphasizing the importance of using uncertainty-aware instructions to improve overall sample efficiency.
@@ 4.1.3 Ablation Study on Uncertainty-aware Instructions
Figure 4: Ablation study on uncertainty-aware instructions. It shows that two types of uncertainty-aware instructions by the teacher both lead to improved sample efficiency for the student agent.
@@
The result of the ablation study is shown in Figure 4. It is shown that utilizing uncertainty-aware instructions improves the overall sample efficiency compared to using deterministic ones. Moreover, there is no significant disparity in performance between the two approaches for generating uncertainty-aware instructions. The first approach is simpler to implement in practical scenarios but consumes more computational resources due to multiple queries to LLMs, particularly when the observation space is large. The second approach necessitates access to logits, making it applicable only to open-source LLMs.
@@ 4.2.1 Experimental Setting
We select the Vicuna-7b model (Chiang et al., 2023) as the LLM used in LLM4Teach, following a similar prompt design as in previous experiments on Minigrid. We utilize visual observations captured by the on-board camera as input queries for the LLM. To enable the LLM-based teacher agent to comprehend these visual inputs, we utilize a pre-configured translator in the same way as in Hu et al. (2024) to generate natural language descriptions which list identified objects in the visual inputs.
@@ 5 Concluding Remarks
Both RL and LLMs have limitations in handling complex sequential decision-making problems. RL often lacks sample efficiency and incurs high exploration costs, while LLMs are prone to decision errors and have high deployment costs. Combining LLMs with RL to overcome these limitations is a natural idea, but creating an effective interface between them poses challenges. LLMs utilize texts as input and output, making them suitable for providing high-level instructions, whereas RL operates at a much lower fine-grained level and uses numerical vectors instead of texts as inputs and outputs.
Here we present LLM4Teach, a novel framework that combines LLMs and RL for embodied sequential decision-making tasks. Our approach leverages the reasoning capabilities of LLMs to develop a highly capable RL-based student agent. In particular, we use the LLM to provide high-level suggestions on available options for policy training of the student agent. Extensive experiments demonstrate that our student agent outperforms all RL baselines in sample efficiency. Meanwhile, it achieves superior performance to LLM soly in terms of task completion success rate with much fewer computational resources during online testing. For instance, in MiniGrid experiments, the student agent’s model size is 24K compared to LLM’s 130B. Similarly, in Habitat experiments, the student agent’s model size is 10M while an LLM’s is 7B.
@@ Acknowledgments
Contribution Statement
Z. Zhou, B. Hu, C. Zhao and P. Zhang contribute equally. B. Liu is project lead.